OpenStudy (anonymous):

Multiply: 4a/5b x 15b/16a x being a multiplication sign, not a variable.

OpenStudy (anonymous):

3/4

OpenStudy (tkhunny):

\(\dfrac{4a}{5b}\cdot\dfrac{15b}{16a} =\dfrac{4}{16}\dfrac{15}{5}\dfrac{a}{a}\dfrac{b}{b}\)